40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 48     Entry time: Thu Feb 4 20:13:53 2010
Author: Frank 
Type: Laser 
Category: Laser 
Subject: output coupler to acav changed 

exchanged the old mirror (T330-HR, T331-AR) by a simple Y1-1025-45P to get more power.

measured laser power : 7.17W
downstream of the new output coupler : 134.6mW

added waveplates & pbs to make the power adjustable. current power through the EOM is 8mW which gives about 4.33V on the RF-PD (Thorlabs PDA10CS, 0dB-setting, 17MHz)

ELOG V3.1.3-